OpenStudy (anonymous):

Subtract. Express your answer in simplest terms. 3 3/4 + 2 1/2

OpenStudy (anonymous):

Is it subtraction or addition?

OpenStudy (anonymous):

addition

OpenStudy (alexandervonhumboldt2):

lol

OpenStudy (anonymous):

Well then first you can split the whole numbers from the fractions so \[\large \sf (3+2)+\huge (\large\frac{3}{4}+\frac{1}{2} \huge )\]

OpenStudy (anonymous):

do we have to find the LCD

OpenStudy (anonymous):

For 3/4 and 1/2, yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i the lcd 4

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

5 4/4

OpenStudy (anonymous):

that is wat i think

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

sorry, but it would be 5/4 = 4/4 + 1/4